HI I Need help here is the problem.
11 over 6p + -7 over 4p

Found 2 solutions by unlockmath, Lina77:
Answer by unlockmath(1688)

Hello,
Let's write it as:
11/6p-7/4p When dealing with fractions that are added or subtracted we need the same denominator. Multiply the first fraction by 2/2 and the second by 3/3 gives us:
22/12p-21/12p Now we can subtract these to get:
1/12p That's the answer.
Make sense?
RJ
Check out my book at www.math-unlock.com

Answer by Lina77(1)

11/6p + -7/4p
Find LCD: LCD is 12p
to get LCD (11/6p * 2/2)+ (-7/4p * 3/3)
Now that you have a common denominator, you can add this fraction properly.
22/12p + -21/12p =
22-21/12p =
Answer: 1/12p